# Evaluating Modern Approaches in 3D Scene Reconstruction: NeRF vs Gaussian-Based Methods

### v1

@@ I Introduction
Recent advancements in Simultaneous Localization and Mapping (SLAM) have significantly leveraged deep learning to enhance the robustness and accuracy of 3D reconstruction and camera tracking. The use of neural implicit representations, as demonstrated by NICE-SLAM [1], allows for detailed reconstructions of large-scale indoor scenes through a hierarchical grid-based encoding that efficiently manages local updates and optimizes scene representations within viewing frustums [2, 1, 3, 4, 5, 6, 7, 8]. Additionally, the comprehensive evaluation in [9] shows that the proposed method outperforms other state-of-the-art approaches, providing a novel solution for automated GI tract segmentation through the integration of specialized architectures.
@@
Comparative analyses between these innovative methods and traditional SLAM systems reveal significant performance improvements. Techniques like GO-SLAM [10] not only outperform older systems like ORB-SLAM2 and DROID-SLAM in tracking accuracy but also in the fidelity of 3D reconstructions [10, 12], especially under challenging conditions that typically result in error accumulation and inconsistencies in traditional methods. Ultimately, this study aims to bridge the gap between theoretical advancements and practical applications, offering insights that could steer future developments in the field of 3D scene reconstruction.
@@ II-B1 NeRF Based
NICE-SLAM [1] employs a structured approach to 3D scene reconstruction, leveraging multi-level voxel grids to enhance detail capture and scalability. This hierarchical grid structure is pivotal in addressing common issues such as the over-smoothing of scene features. By updating only the grid features that are visible, NICE-SLAM significantly improves optimization precision and operational efficiency, contrasting sharply with methods like iMAP which rely on global updates and may suffer from inefficiency and error propagation. The algorithm typically utilizes a set of predefined parameters for constructing the voxel grids [15, 16]. These include setting the grid resolutions at different hierarchy levels to balance detail and computational efficiency. For instance, coarse grids capture basic structural outlines, while finer grids focus on detailed textures and objects.
However, NICE-SLAM is not without its drawbacks. Its predictive performance is bounded by the resolution of the coarsest grid, which may limit its applicability in scenarios requiring high precision over large scales [17, 18].
@@
Point-SLAM [19] introduces a dynamic neural point cloud approach that adapts density based on the detailed requirements driven by the data, significantly enhancing memory efficiency. By utilizing per-pixel image gradients, the model smartly modulates point density, concentrating on areas requiring more detail and simplifying regions with less complexity. As exploration progresses, Point-SLAM expands the point cloud, optimizing spatial usage by focusing point density and compressing points in less detailed regions, thereby maintaining computational efficiency in real-time scene reconstruction [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30].
In Point-SLAM, the model’s parameterization is crucial for fine-tuning its dynamic neural point cloud. Typical parameters include the resolution of the point cloud, which adjusts dynamically to match the density required by the scene complexity. The model also leverages parameters related to the gradient thresholds that control point density based on per-pixel image gradients [31]. Additionally, parameters for space optimization ensure efficient memory use by managing how points are distributed and compressed in areas of lesser detail. These parameters are key to maintaining the balance between detail fidelity and computational efficiency in real-time applications.
@@ II-B2 Gaussian Based
The first step in SplaTAM involves data acquisition from RGB-D cameras, where both color and depth information are collected. The depth data is filtered and aligned with the RGB frames using methods from [33], ensuring consistency in representation. Subsequently, the algorithm constructs 3D Gaussians, known as ”splats,” from the incoming depth data. Each splat is represented by its mean, covariance, and color, enabling a rich representation of the scene’s geometry and appearance [34]. For tracking, SplaTAM employs an optimized keyframe selection technique, enabling significant computational savings while maintaining accurate localization. The framework utilizes a particle filter that tracks the agent’s pose by propagating a set of hypotheses through the splat-based model [35, 36, 37, 38, 39]. Each particle incorporates splat information, allowing for multi-hypothesis tracking, which is crucial in dynamic environments. The data association process employs a likelihood-based method, assessing the correspondence between observed features and existing splats to update particle weights [40]. Mapping in SplaTAM is accomplished by integrating new observations into the existing splat representation in real-time using a Kalman filter approach, adjusting their parameters based on the incoming data [41, 42, 43, 44, 45]. The algorithm intelligently manages occlusions and dynamic elements by classifying splats as either static or dynamic, utilizing robust outlier rejection strategies [46] to ensure the integrity of the static map over time.
Extensive evaluations demonstrate that SplaTAM outperforms traditional SLAM systems in terms of mapping fidelity and localization precision, making it suitable for real-time applications in challenging scenarios [32]. Metrics such as geometric consistency and computational efficiency are continuously monitored to ensure robust performance, providing an essential framework for navigating complex and dynamic environments [47]. However, SplaTAM’s computational complexity may hinder real-time performance, particularly in resource-constrained environments [48].

### v2

@@ I Introduction
Recent advancements in Simultaneous Localization and Mapping (SLAM) have significantly leveraged deep learning to enhance the robustness and accuracy of 3D reconstruction and camera tracking. The use of neural implicit representations, as demonstrated by NICE-SLAM [1], allows for detailed reconstructions of large-scale indoor scenes through a hierarchical grid-based encoding that efficiently manages local updates and optimizes scene representations within viewing frustums [2, 1, 3]. Additionally, the comprehensive evaluation in [4] shows that the proposed method outperforms other state-of-the-art approaches, providing a novel solution for automated GI tract segmentation through the integration of specialized architectures.
@@
Comparative analyses between these innovative methods and traditional SLAM systems reveal significant performance improvements. Techniques like GO-SLAM [5] not only outperform older systems like ORB-SLAM2 and DROID-SLAM in tracking accuracy but also in the fidelity of 3D reconstructions [5], especially under challenging conditions that typically result in error accumulation and inconsistencies in traditional methods. Ultimately, this study aims to bridge the gap between theoretical advancements and practical applications, offering insights that could steer future developments in the field of 3D scene reconstruction.
@@ II-B1 NeRF Based
NICE-SLAM [1] employs a structured approach to 3D scene reconstruction, leveraging multi-level voxel grids to enhance detail capture and scalability. This hierarchical grid structure, as illustrated in Fig. 1, is pivotal in addressing common issues such as the over-smoothing of scene features. By updating only the grid features that are visible, NICE-SLAM significantly improves optimization precision and operational efficiency, contrasting sharply with methods like iMAP which rely on global updates and may suffer from inefficiency and error propagation. The algorithm typically utilizes a set of predefined parameters for constructing the voxel grids [9]. These include setting the grid resolutions at different hierarchy levels to balance detail and computational efficiency. For instance, coarse grids capture basic structural outlines, while finer grids focus on detailed textures and objects.
However, NICE-SLAM is not without its drawbacks. Its predictive performance is bounded by the resolution of the coarsest grid, which may limit its applicability in scenarios requiring high precision over large scales [10].
@@
Point-SLAM [11] introduces a dynamic neural point cloud approach that adapts density based on the detailed requirements driven by the data, significantly enhancing memory efficiency. By utilizing per-pixel image gradients, the model smartly modulates point density, concentrating on areas requiring more detail and simplifying regions with less complexity. As exploration progresses, Point-SLAM expands the point cloud, optimizing spatial usage by focusing point density and compressing points in less detailed regions, thereby maintaining computational efficiency in real-time scene reconstruction [12, 13, 14].
In Point-SLAM, the model’s parameterization is crucial for fine-tuning its dynamic neural point cloud. Typical parameters include the resolution of the point cloud, which adjusts dynamically to match the density required by the scene complexity. The model also leverages parameters related to the gradient thresholds that control point density based on per-pixel image gradients. Additionally, parameters for space optimization ensure efficient memory use by managing how points are distributed and compressed in areas of lesser detail. These parameters are key to maintaining the balance between detail fidelity and computational efficiency in real-time applications.
@@ II-B2 Gaussian Based
The first step in SplaTAM involves data acquisition from RGB-D cameras, where both color and depth information are collected. The depth data is filtered and aligned with the RGB frames using methods from [16], ensuring consistency in representation. Subsequently, the algorithm constructs 3D Gaussians, known as ”splats,” from the incoming depth data. Each splat is represented by its mean, covariance, and color, enabling a rich representation of the scene’s geometry and appearance [17]. For tracking, SplaTAM employs an optimized keyframe selection technique, enabling significant computational savings while maintaining accurate localization. The framework utilizes a particle filter that tracks the agent’s pose by propagating a set of hypotheses through the splat-based model. Each particle incorporates splat information, allowing for multi-hypothesis tracking, which is crucial in dynamic environments. The data association process employs a likelihood-based method, assessing the correspondence between observed features and existing splats to update particle weights [18]. Mapping in SplaTAM is accomplished by integrating new observations into the existing splat representation in real-time using a Kalman filter approach, adjusting their parameters based on the incoming data. The algorithm intelligently manages occlusions and dynamic elements by classifying splats as either static or dynamic, utilizing robust outlier rejection strategies [19] to ensure the integrity of the static map over time.
Extensive evaluations demonstrate that SplaTAM outperforms traditional SLAM systems in terms of mapping fidelity and localization precision, making it suitable for real-time applications in challenging scenarios [15]. Metrics such as geometric consistency and computational efficiency are continuously monitored to ensure robust performance, providing an essential framework for navigating complex and dynamic environments. However, SplaTAM’s computational complexity may hinder real-time performance, particularly in resource-constrained environments.